INVESTIGATION OF VAN DER WAALS COMPLEXES IN A FREE EXPANSION OF C_{2}H_{2}/X (X=NOBLE GAS) USING CW CAVITY RING-DOWN SPECTROSCOPY IN THE OVERTONE RANGE

\underline{C. LAUZIN}, K. DIDRICHE, T. FOLDES, M. HERMAN, Service de Chimie quantique et Photophysique CP160/09, Faculté des Sciences, Université Libre de Bruxelles (U.L.B.), Av. Roosevelt, 50, B-1050, Bruxelles, Belgium.

CW cavity ring down spectroscopy around 1.5 μm was used to probe a free supersonic expansion of mixed acetylene and noble gas (He, Ne, Ar and Kr). Pure C_{2}H_{2} and mixed complexes are observed at sub-Doppler resolution around the 2CH acetylenic band. The effects of the different noble gases on the complexation are discussed. The updated or new spectral analysis of the spectra will be presented.